What is the mass in grams of oxygen gas, O2 is needed to complete the combustion of 6 L of methane, CH4? Assume that the pressure and temperature remain constant. (be sure to write the balanced chemical equation)


1
Expert's answer
2021-11-13T01:33:45-0500

CH4 + 2O2 → CO2 + 2H2O

One mole of any gas at STP occupy volume of 22.4 L.

Proportion:

22.4 L – 1 mol

6 L – x

"x = \frac{6 \times 1}{22.4} = 0.2678 \;mol"

According to the reaction:

n(O2) = 2n(CH4) "= 2 \times 0.2678 = 0.5356 \;mol"

M(O2) = 32 g/mol

m(O2) "= 0.5356 \times 32 = 17.14 \;g"

Answer: 17.14 g